According to the following sequence how to find the nth term
1, 2, 3, 6, 11, 20, 37, 68, …

Answer by greenestamps(13200)



TRIbonacci numbers -- like Fibonacci numbers except each term is the sum of the preceding THREE terms:
first three numbers 1, 2, 3;
each subsequent term is the sum of the preceding three terms

1
2
3
1+2+3=6
2+3+6=11
3+6+11=20
6+11+20=37
11+20+37=68
...

You can use the recursive definition to find as many terms as you want.

I don't believe there is an explicit formula for the n-th term.